# Interpretable and Equation-Free Response Theory for Complex Systems

### v2

@@ II.2 Second order response
where we have used the expression for $\nu^{(1)}$ above. Note that the $\Theta$’s ensure the correct time ordering of the perturbation.
@@
Note again that our summation excludes the term corresponding to the invariant measure of the systen. The results above can easily be extended at all orders of perturbations, see Appendix A. This implies that the full nonlinear time-dependent response can be obtained from the knowledge of $m$, $\mathcal{M}$, and $f$ for any observable of the system. Assuming $f=1$ and taking the $n\rightarrow\infty$ limit, the results presented in [41] are easily recovered.
III Response theory for Correlations
@@ IV A Simple Example
where $\mathrm{d}W_{1}$ and $\mathrm{d}W_{2}$ are increments of independent Wiener processes, and
$V(x,y)$ $=3\exp(-x^{2}-(y-1/3)^{2})-3\exp(-x^{2}-(y-5/3)^{2})$
@@
The example is taken from [49] with a small modification. The potential is depicted in Fig. 1a and features three local minima, located at $(x_{1},y_{1})\approx(0,1.55)$, $(x_{2},y_{2})=(-1,0)$, and $(x_{3},y_{3})=(1,0)$. The potential we consider here corresponds almost exactly to a case study presented in [49]. We have included an additional term $-y$ in the definition of the potential in order to lower the local minimum in $(x_{1},y_{1})$ to a value closer to the absolute minimum that is realized at $(x_{2},y_{2})$ and $(x_{3},y_{3})$.
@@
Figure 1: (a) Potential function $V(x,y)$ with approximate indication of the three quasi-invariant regions surrounding the minina of the $V$.(b) Invariant Measure $\rho_{0}\propto\exp(-2V(x,y)/\sigma^{2})$. (c). First subdominant Kolmogorov mode, $\lambda_{1}=0.9916(2)$. (d) First subdominant mode of the Perron-Frobenius Operator. (e) Second subdominant Kolmogorov model, $\lambda_{2}=0.9655(2)$. (f) Second subdominant mode of the Perron-Frobenius Operator.
@@
The Fokker-Planck equation associated with the previous Langevin equation reads as
$\partial_{t}\rho=\mathcal{L}_{0}\rho=\partial_{x}(\partial_{x}V\rho)+\partial_{y}(\partial_{y}V\rho)+\frac{\sigma^{2}}{2}(\partial_{x}^{2}+\partial_{y}^{2})\rho.$ (26)
@@
where $\mathcal{L}_{0}$ is the Fokker-Planck operator. We have that $P^{\tau,0}=\exp(\mathcal{L}_{0}\tau)$ is the associated Perron-Frobenius operator describing the evolution of measures for a time $\tau$. We refer to the The unique invariant measure in depicted in Fig. 1b, obeys $\mathcal{L}_{0}\rho_{0}=0$, and is given by $\rho_{0}=Z^{-1}\exp(-2V/\sigma^{2})$, where $Z$ is the normalization factor. In what follows, we assume $\sigma=0.7$. Our numerical simulations are performed using the standard Euler-Maruyama scheme [61] with $dt=0.05$. We sample our output every 20 times steps, so that our reference time scale is $\tau=1$. The dynamics given by Eqs. 23-26 describes an equilibrium system.
@@
To approximate $P^{\tau,0}$, we discretize the phase space of the system into $N$ disjoint subsets $\{B_{1},B_{2},\dots,B_{N}\}$, forming a partition of the space. The Ulam transfer operator $P^{\tau,0}_{\{N\},ij}$, a finite-dimensional stochastic matrix describes the probability that the orbit of the system is at time $t+\tau$ in the subset $B_{i}$ is at time $t$ it was in the subset $B_{j}$. In our case, the $B_{j}$’s are given by the 625 cubes with side 0.2 centred around the origin. Whilst in principle one would need to cover the entire $\mathbb{R}^{2}$ the confining potential makes all regions beyond those we consider here entirely irrelevant unless one consider extremely long time scales.
@@
Following [49], we populate each cube with 1000 ensemble members distributed uniformly according to the Lebesgue measure. Each member evolves for $\tau=1$ time unit. We then construct an estimate of $P^{\tau,0}_{\{N\}}$ by counting the transitions. We repeat the operation 20 times and by averaging we obtain our best estimate of $P^{\tau,0}_{\{N\}}$, which constitutes our reference discretized stochastic matrix, so that $\mathcal{M}=P^{\tau,0}_{\{N\}}$. The solution to $\mathcal{M}\nu_{inv}=\nu_{inv}$ gives an extremely good approximation in the gridded domain defined by the $B_{j}$’s of the true invariant measure $\rho_{0}$. The discretized system obeys detailed balance, as expected from the nature of the original continuum system. The corresponding left eigenvector, which is the first right eigenvector of the Koopman operator $\mathcal{M}^{T}$, is constant everywhere. We obtain 624 additional eigenvalue-eigenvector pairs for the matrix $\mathcal{M}$. Since the system obeys detailed balance, we would expect all of them to be real, but, because of insufficient sampling and numerical precision, this constraint is not necessarily obeyed for the rapidly decaying ones.
@@
The two subdominant eigenvectors are depicted in Fig. 1d and Fig. 1f. The first one describes the transitions between the neighbourhoods of $(x_{2},y_{2})$ and $(x_{3},y_{3})$. The second one describes the transitions between the neighbourhoods of $(x_{1},y_{1})$ and either $(x_{2},y_{2})$ or $(x_{3},y_{3})$. The corresponding subdominant eigenvectors of the Koopman operator are depicted in Fig. 1c and Fig. 1e. There is a very large spectral gap between the three dominant modes (corresponding to $\lambda=1$, $\lambda\approx 0.9917(2)$, and $\lambda\approx 0.9655(2)$) and the rest of the spectral components ($\lambda_{4}\approx 0.050(2))$, which indicates that the system can safely undergo a model reduction procedure. Note that the uncertainty on the least significant digit has been estimated using 10 separate integrations. The model reduction could be algebraically achieved by substituting $\mathcal{K}\rightarrow\sum_{i=1}^{3}\lambda_{i}\Pi_{i}$, and $\mathcal{M}\rightarrow\sum_{i=1}^{3}\lambda_{i}\Pi_{i}^{T}$. Additionally, the level sets of the two subdominant Koopman modes can be used to identify the three metastable regimes of the system, which are indicated as (1), (2) and (3) in Fig. 1a and correspond to the basins of attraction of the three local minima in the case the stochastic forcing is switched off.
@@
We use such a geometrical partition of the phase space to introduce a separate reduced-order representation of the dynamics, whereby the phase space of the system is partitioned into three states, corresponding to the regions (1), (2), and (3). In the simple case described here, this corresponds to MSM. This amounts to neglecting entirely the intrawell dynamics, which, in the Ulam description above, is mostly captured by the Koopman modes with index $\geq 4$. We estimate the reduced Markov model (RMM) discretized transfer operator $\tilde{M}$ by performing a single run lasting $10^{7}$ time units (after disregarding a small transient) and counting the transitions between the 3 states described above. The obtained estimates for $\tilde{M}$ and its eigenvectors are given below, where all the numbers indicated below have an approximate uncertainty of 2 in the least significant we have written out.
@@
Also this minimal model obeys detailed balance and the two nontrivial eigenvalues $\lambda_{2}=0.9904(2)$ and $\lambda_{3}=0.9606(2)$ correspond closely to the first two subdominant eigenvalues obtained using the Ulam method, and the corresponding eigenvectors provide a coarse grained version of the figures provided in Fig. 1d and Fig. 1f.
IV.1 Applying Extra Forcings
@@
$\langle\Phi\rangle(t)=\langle\Phi\rangle_{0}+\epsilon_{1}\int_{-\infty}^{\infty}\mathrm{d}t_{1}G^{(1)}_{x,\Psi}(t-t_{1})f_{1}(t_{1})+\epsilon_{2}\int_{-\infty}^{\infty}\mathrm{d}t_{1}G^{(1)}_{y,\Psi}(t-t_{1})f_{2}(t_{1})+h.o.t.$ (27)
where $\langle\Phi\rangle_{0}=\int\mathrm{d}x\mathrm{d}y\rho_{0}(x,y)\Psi(x,y)$ is the expectation value of $\Psi$ in the unperturbed asymptotic state, whilst
@@
IV.1.1 Linear Response - Observables
@@
We choose as observables $\Psi_{1}=x$ and $\Psi_{2}=y$. We first note that because of the symmetry of the system $G^{(1)}_{x,y}(\tau)=G^{(1)}_{y,x}(\tau)=0$. We then focus on the case where we force the system in the $j$ direction and use $j$ as observable, $j=x,y$. We then estimate $G^{(1)}_{x/y,x/y}$ through formula 28 by collecting statistics for 500 independent ensemble runs of the unpertubed system each lasting $10^{6}$ time units. Note that this 50 times as many data as those used for constructing the Markov chains. From the knowledge of $m_{x},$, $m_{y}$, it is straightforward to compute the Green’s functions for $\Psi_{1}=x$ and $\Psi_{2}=y$ using Eq. II.1.
The estimates we obtain for the Green’s functions of interest are shown in Fig. 2. When considering RMM, we clearly see that $\mathcal{G}^{(1)}_{m_{x},x}(\tau)\propto\lambda_{2}^{\tau}$ and $\mathcal{G}^{(1)}_{m_{y},y}(\tau)\propto\lambda_{3}^{\tau}$, which as a result of the choice of forcing/observable pair, only one Koopman mode is retained in the expansion given in Eqs. 7- 8. Indeed, it is clear that if we force along the $x$-direction and choose $x$ as observable, the second Koopman mode (or only the second Perron-Frobenium mode) is the only one that matters. Similarly, the third mode is the only one retained in the spectral expansion of the Green’s function when forcing along $y$-direction and choosing $y$ as observable.
@@
It is apparent that the estimates of the two Green’s functions obtained using the FDT are relatively noisy, despite the use of a many times more data than those used for constructing the Markov chains. A side remark is that, by construction, $\lim_{t\rightarrow 0^{+}}\mathcal{G}^{(1)}_{m_{x},x}(t)=\lim_{t\rightarrow 0^{+}}\mathcal{G}^{(1)}_{m_{y},y}(t)=1$. Nonetheless, the function collapses to much lower values within one time unit because of the very rapid decay of correlation due to the rapidly decaying Kolmogorov modes of the continuum system. As soon as we consider $t\geq 1$, a rather good agreement is found with the estimates obtained via Markov chains.
IV.1.2 Linear Response - Correlations
@@
In the case of the Markov chain, we have used the expression of correlations and the linear response formulas to static forcings provided in Sect. III.
We obtain that - see Figs. 3a), 3a)c), and3e) - there is very good agreement between the correlations computed on the full system and those estimated via Markov chains. Comparing with Fig. 2, it is apparent that $C_{\tau}(x,x)|_{\epsilon_{2}=\epsilon_{1}=0}\approx\exp(\lambda_{2}t)$ and $C_{\tau}(y,y)|_{\epsilon_{2}=\epsilon_{1}=0}\approx\exp(\lambda_{3}t)$.
@@
Figure 3: (a) Estimate of $C_{\tau}(x,x)\rangle$ via direct numerical simulations (DNS), RMM, and Ulam method. Red lines: Reference state. Black lines: sensitivity with respect to $\epsilon_{2}$. (b) Decomposition of the linear response in (a) in the four terms discussed in Eq. III (RMM). (c) Same as (a), but for $C_{\tau}(y,x)$ and its sensitivity with respect to $\epsilon_{1}$. (d) Same as (b), in reference to the linear response shown in (c). (e) Same as (a), but for $C_{\tau}(y,y)$ and its sensitivity with respect to $\epsilon_{2}$. (f). Same as (b), in reference to the linear response shown in (e).
Similarly, a good agreement is found when considering the sensitivities with respect to $\epsilon_{1}$ and $\epsilon_{2}$. In all cases it is apparent that a little fraction of the signal is lost when performing the coarse graining. Nonetheless, the good performance obtained even in the case of the 3-state system indicates the effectiveness of the reduced order modelling strategy. Using the decomposition presented in Eq. III, we are able to separate the change in the correlation function in four components, which are associated with fundamentally different dynamical processes. Two terms come from the sensitivity in the expectation value of the two observables we are considering. An additional term - indicated by $\delta^{(1)}_{b,\epsilon}$ comes from the change in the expectation value of the lagged product of the observable due exclusively to the change in the measure (where instead the evolution occurs according to the unperturbed dynamics). The most interesting term is undoubtedly $\delta^{(1)}_{a,\epsilon}$, which measures the impact of the change in the dynamics occurring up to the considered time lag. Indeed, this terms vanishes as $\tau\rightarrow 0$ and is, as already observed in [48], a specific element of response formulas for correlations. The interplay between the two terms $\delta^{(1)}_{a,\epsilon}$ and $\delta^{(1)}_{b,\epsilon}$ is nontrivial.
IV.1.3 Nonlinear Response

### v3

@@ 2.2 Second order response
where we have used the expression for $\nu^{(1)}$ above. Note that the $\Theta$’s ensure a correctly time-ordered consideration of the acting perturbation.
@@
Note again that our summation excludes the term corresponding to the invariant measure of the system. The results above can easily be extended at all orders of perturbations, see Appendix A. This implies that the full nonlinear time-dependent response can be obtained from the knowledge of $m$, $\mathcal{M}$, and $f$ for any observable of the system. Assuming $f=1$ and taking the $n\rightarrow\infty$ limit, the results presented in [59] can be easily recovered.
3 Response theory for Correlations
@@ 4 A Simple Example
where $\mathrm{d}W_{1}$ and $\mathrm{d}W_{2}$ are increments of independent Wiener processes, and the drift is defined as minus the gradient of the following potential:
$V(\mathbf{x})=V(x,y)$ $=3\exp(-x^{2}-(y-1/3)^{2})-3\exp(-x^{2}-(y-5/3)^{2})$
@@
Given the choice of the drift and of the noise law, we are considering here an equilibrium system obeying detailed balance [7]. The potential is depicted in Fig. 1a and features three local minima, located at $(x_{1},y_{1})\approx(0,1.55)$, $(x_{2},y_{2})=(-1,0)$, and $(x_{3},y_{3})=(1,0)$. The potential we consider here corresponds almost exactly to a case study presented in [68]. We have included an additional term $-y$ in the definition of the potential in order to reduce the value of $V$ at the local minimum in $(x_{1},y_{1})$ to a value closer to the absolute minimum that is realized at $(x_{2},y_{2})$ and $(x_{3},y_{3})$.
@@
Figure 1: (a) Potential function $V(x,y)$ with approximate indication of the three quasi-invariant regions surrounding the minima of the $V$.(b) Invariant Measure $\rho_{0}\propto\exp(-2V(x,y)/\sigma^{2})$. (c). First subdominant Koopman mode, $\lambda_{1}=0.9916$. (d) First subdominant mode of the Perron-Frobenius Operator. (e) Second subdominant Koopman model $\lambda_{2}=0.9655$. (f) Second subdominant mode of the Perron-Frobenius Operator.
@@
The Fokker-Planck equation associated with the previous Langevin equation reads as [83, 7]:
$\partial_{t}\rho(\mathbf{x},t)=\mathcal{L}_{0}\rho(\mathbf{x},t)=\partial_{x}(\partial_{x}V(\mathbf{x})\rho(\mathbf{x},t))+\partial_{y}(\partial_{y}V(\mathbf{x})\rho(\mathbf{x},t))+\frac{\sigma^{2}}{2}(\partial_{x}^{2}+\partial_{y}^{2})\rho(\mathbf{x},t)$ (26)
The solution can be expressed as $\rho(\mathbf{x},t)=\exp(\mathcal{L}_{0}t)\rho_{in}(\mathbf{x})$ where $\rho_{in}(\mathbf{x})$ is the initial condition and $\exp(\mathcal{L}_{0}t)$ is the strongly continuous semigroup generated by $\mathcal{L}_{0}$, which is densely defined in a suitable Banach space [7, 16]. The Perron-Frobenius operator $P^{t,0}=\exp(\mathcal{L}_{0}t)$ describes the evolution of measures for a time $t$. Since we have a confining potential for the drift and non-degenerate noise, we have a unique invariant measure that is absolutely continuous with respect to Lebesgue, so that it can be written as $\rho_{0}\mathrm{d}\mathbf{x}$, where $\mathcal{L}_{0}\rho_{0}=0$ [7, 84]. We obtain $\rho_{0}(\mathbf{x})=Z^{-1}\exp(-2V(\mathbf{x})/\sigma^{2})$, where $Z$ is the normalisation factor ensuring that $\int\mathrm{d}\mathbf{x}\rho_{0}(\mathbf{x})=1$. As a result, the natural functional space for our problem is given by $L^{2}_{\rho_{0}}$. The Koopman operator $K^{t,0}$ on measurable observables and is defined as the adjoint of the Perron-Frobenius operator: $\int K^{t,0}\Psi(\mathbf{x})\rho_{in}(\mathbf{x})\stackrel{{\scriptstyle\text{def}}}{{=}}\int\mathrm{d}\mathbf{x}\Psi(\mathbf{x})P^{t,0}\rho_{in}(\mathbf{x})$, where $K^{t,0}=(P^{t,0})^{T}=\exp(\mathcal{L}^{T}_{0}t)$ and $\mathcal{L}^{T}_{0}\Psi(\mathbf{x})=-\partial_{x}V(\mathbf{x})\partial_{x}\Psi(\mathbf{x})-\partial_{y}V(\mathbf{x})\partial_{y}\Psi(\mathbf{x})+\frac{\sigma^{2}}{2}(\partial^{2}_{x}+\partial^{2}_{y})\Psi(\mathbf{x})$.
@@
In what follows, we assume $\sigma=0.7$. Our numerical simulations are performed using the standard Euler-Maruyama scheme [85] with $dt=0.05$. We sample our output every 20 times steps, so that our reference time scale is $\tau=1$. The dynamics given by Eqs. 23-26 describes an equilibrium system.
@@
To approximate $P^{\tau,0}$, we discretize the phase space of the system into $N$ disjoint subsets $\{B_{1},B_{2},\dots,B_{N}\}$, forming a partition of the space. The Ulam transfer operator $P^{\tau,0}_{\{N\},ij}$, a finite-dimensional stochastic matrix describes the probability that the orbit of the system is at time $t+\tau$ in the subset $B_{i}$ is at time $t$ it was in the subset $B_{j}$. In our case, the $B_{j}$’s are given by the 625 squares with side 0.2 centred around the origin. Whilst in principle one would need to cover the entire $\mathbb{R}^{2}$ the confining potential makes all regions beyond those we consider here entirely irrelevant unless one consider extremely long time scales.
@@
Following [68], we populate each cube with 1000 ensemble members distributed uniformly according to the Lebesgue measure. Each member evolves for $\tau=1$ time unit. We then construct an estimate of $P^{\tau,0}_{\{N\}}$ by counting the transitions. We repeat the operation 20 times and by averaging we obtain our best estimate of $P^{\tau,0}_{\{N\}}$, which constitutes our reference discretized stochastic matrix, so that $\mathcal{M}=P^{\tau,0}_{\{N\}}$. The solution to $\mathcal{M}\nu_{inv}=\nu_{inv}$ gives an accurate approximation in the gridded domain defined by the $B_{j}$’s of the true invariant measure $\rho_{0}$.
It is possible to filter out the dynamics associated with spurious non-equilibrium modes due to the finite sampling by proceeding as follows. First, we compute the matrix $\mathcal{N}_{ij}=\mathcal{M}_{ji}\frac{\nu_{inv,i}}{\nu_{inv,j}}$. Them we update our estimate of $\mathcal{M}$ as follows: ${\mathcal{M}}\rightarrow\frac{\mathcal{M}+\mathcal{N}}{2}$, so that detailed balance is enforced in the finite-state representation of the system [^2] If $\mathcal{M}$ obeys detailed balance, the symmetric matrix $h=D\mathcal{M}D^{-1}$, where $D=diag(\nu_{inv,1},\ldots,\nu_{inv,N})$ has the same eigenvalues as those of $\mathcal{M}$. Estimating the eigenvalues via $h$ could prove helpful for large $N$ thanks to its numerically convenient symmetry property..
@@
Whilst the first eigenvector of $\mathcal{M}$ is shown in Fig. 1b, the first right eigenvector of the Koopman operator $\mathcal{M}^{T}$, is constant everywhere. We obtain 624 additional eigenvalue-eigenvector pairs for the matrix $\mathcal{M}$. The two subdominant eigenvectors are depicted in Fig. 1d and Fig. 1f. The first one describes the transitions between the neighbourhoods of $(x_{2},y_{2})$ and $(x_{3},y_{3})$. The second one describes the transitions between the neighbourhoods of $(x_{1},y_{1})$ and either $(x_{2},y_{2})$ or $(x_{3},y_{3})$. The corresponding subdominant eigenvectors of the Koopman operator are depicted in Fig. 1c and Fig. 1e. There is a very large spectral gap between the three dominant modes (corresponding to $\lambda=1$, $\lambda\approx 0.9917(1)$, $\lambda\approx 0.9655(1)$ the rest of the spectral components ($\lambda_{4}\approx 0.050(1))$, which indicates that the system can safely undergo a model reduction procedure. This could be algebraically achieved by substituting $\mathcal{K}\rightarrow\sum_{i=1}^{3}\lambda_{i}\Pi_{i}$, and $\mathcal{M}\rightarrow\sum_{i=1}^{3}\lambda_{i}\Pi_{i}^{T}$. Additionally, the level sets of the two subdominant Koopman modes can be used to identify the three metastable regimes of the system, which are indicated as (1), (2) and (3) in Fig. 1a and correspond to the basins of attraction of the three local minima in the case the stochastic forcing is switched off. Performing k-means clustering [37, 38] of the data choosing 3 as number of clusters leads to the same partition of the data space, whereby the three regions (1), (2), (3) are the Voronoï cells [36] associated with the cluster centers.
@@
We use such a geometrical partition of the phase space to introduce a separate reduced-order representation of the dynamics, whereby the phase space of the system is partitioned into three states, corresponding to the regions (1), (2), and (3). In the simple case described here, this corresponds to MSM. This amounts to neglecting entirely the intrawell dynamics, which, in the Ulam description above, is mostly captured by the Koopman modes with index $\geq 4$. We estimate the reduced Markov model (RMM) discretized transfer operator ${M}$ by performing a single run lasting $10^{7}$ time units (after disregarding a small transient) and counting the transitions between the 3 states described above. We then enforce detailed balance as above. The obtained estimates for ${M}$ and its eigenvectors are given below, where all the numbers indicated below have an approximate uncertainty of 1 in the last digit we have written out.
@@
The two nontrivial eigenvalues $\lambda_{2}=0.9904(1)$ and $\lambda_{3}=0.9606(1)$ correspond closely to the first two subdominant eigenvalues obtained using the Ulam method, and the corresponding eigenvectors provide a coarse grained version of the figures provided in Fig. 1d and Fig. 1f.
4.1 Applying Extra Forcings
@@
$\langle\Psi\rangle(t)=\langle\Psi\rangle_{0}+\epsilon_{1}\int_{-\infty}^{\infty}\mathrm{d}t_{1}G^{(1)}_{x,\Psi}(t-t_{1})f_{1}(t_{1})+\epsilon_{2}\int_{-\infty}^{\infty}\mathrm{d}t_{1}G^{(1)}_{y,\Psi}(t-t_{1})f_{2}(t_{1})+h.o.t.$ (27)
where $\langle\Psi\rangle_{0}=\int\mathrm{d}x\mathrm{d}y\rho_{0}(x,y)\Psi(x,y)$ is the expectation value of $\Psi$ in the unperturbed state, whilst
@@
4.1.1 Linear Response - Observables
@@
We choose as observables $\Psi_{1}=x$ and $\Psi_{2}=y$. We first note that because of the symmetry of the system $G^{(1)}_{x,y}(\tau)=G^{(1)}_{y,x}(\tau)=0$. We then focus on the case where we force the system in the $j$ direction and use $j$ as observable, $j=x,y$. We then estimate $G^{(1)}_{x/y,x/y}$ through formula 28 by collecting statistics for 500 independent ensemble runs of the unpertubed system each lasting $10^{6}$ time units. Note that this 50 times as many data as those used for constructing the Markov chains. From the knowledge of $m_{x},$, $m_{y}$, it is straightforward to compute the Green’s functions for $\Psi_{1}=x$ and $\Psi_{2}=y$ using Eq. 2.1.
The estimates we obtain for the Green’s functions of interest are shown in Fig. 2. When considering RMM, we clearly see that $\mathcal{G}^{(1)}_{m_{x},x}(\tau)\propto\lambda_{2}^{\tau}$ and $\mathcal{G}^{(1)}_{m_{y},y}(\tau)\propto\lambda_{3}^{\tau}$: as a result of the choice of forcing/observable pair, only one Koopman mode is retained in the expansion given in Eqs. 7- 8. Indeed, it is clear that if we force along the $x$-direction and choose $x$ as observable, the second Koopman mode (or only the second Perron-Frobenium mode) is the only one that given a nonnvanishing contribution. Similarly, the third mode is the only one retained in the spectral expansion of the Green’s function when forcing along $y$-direction and choosing $y$ as observable.
@@
It is apparent that the estimates of the two Green’s functions obtained using the FDT are relatively noisy, despite the use of a much larger dataset than what has been used for constructing the Gree’s function via Markov chains. A side remark is that, by construction, $\lim_{t\rightarrow 0^{+}}\mathcal{G}^{(1)}_{m_{x},x}(t)=\lim_{t\rightarrow 0^{+}}\mathcal{G}^{(1)}_{m_{y},y}(t)=1$. Nonetheless, the function collapses to much lower values within one time unit because of the very rapid decay of correlation due to the rapidly decaying Kolmogorov modes of the continuum system. As soon as we consider $t\geq 1$, a rather good agreement is found with the estimates obtained via Markov chains.
4.1.2 Linear Response - Correlations
@@
In the case of the Markov chain, we have used the expression of correlations and the linear response formulas to static forcings provided in Sect. 3.
We obtain that - see Figs. 3a), 3a)c), and 3e) - there is a very good agreement between the correlations computed on the full system and those estimated via Markov chains. Comparing with Fig. 2, it is apparent that $C_{\tau}(x,x)|_{\epsilon_{2}=\epsilon_{1}=0}\propto\exp(\lambda_{2}t)$ and $C_{\tau}(y,y)|_{\epsilon_{2}=\epsilon_{1}=0}\propto\exp(\lambda_{3}t)$.
@@
Figure 3: (a) Estimate of $C_{\tau}(x,x)\rangle$ via direct numerical simulations (DNS), RMM, and Ulam method. Red lines: Reference state. Black lines: sensitivity with respect to $\epsilon_{2}$. (b) Decomposition of the linear response in (a) in the four terms discussed in Eq. 3 (RMM). (c) Same as (a), but for $C_{\tau}(y,x)$ and its sensitivity with respect to $\epsilon_{1}$. (d) Same as (b), in reference to the linear response shown in (c). (e) Same as (a), but for $C_{\tau}(y,y)$ and its sensitivity with respect to $\epsilon_{2}$. (f). Same as (b), in reference to the linear response shown in (e).
Similarly, a good agreement is found when considering the sensitivities with respect to $\epsilon_{1}$ and $\epsilon_{2}$. In all cases it is apparent that a little fraction of the signal is lost when performing the coarse graining. The good performance obtained even in the case of the 3-state system indicates the effectiveness of the reduced order modelling strategy. Using the decomposition presented in Eq. 3, we are able to separate the change in the correlation function in four components, which are associated with fundamentally different dynamical processes. Two terms come from the sensitivity in the expectation value of the two observables we are considering. An additional term - indicated by $\delta^{(1)}_{b,\epsilon}$ comes from the change in the expectation value of the lagged product of the observable due exclusively to the change in the measure (where instead the evolution occurs according to the unperturbed dynamics). The most interesting term is undoubtedly $\delta^{(1)}_{a,\epsilon}$, which measures the impact of the change in the dynamics occurring up to the considered time lag. Indeed, this terms vanishes as $\tau\rightarrow 0$ and is, as already observed in [67], a specific element of response formulas for correlations. The interplay between the two terms $\delta^{(1)}_{a,\epsilon}$ and $\delta^{(1)}_{b,\epsilon}$ is nontrivial.
4.1.3 Nonlinear Response